OpenStudy (anonymous):

how to find the zeros of \[x ^{3}\] + x + 10

OpenStudy (nikvist):

\[x^3+x+10=0\] \[(x+2)\cdot(x^2-2x+5)=0\] \[(x+2)\cdot((x-1)^2+4)=0\] \[x_1=-2\quad,\quad x_{2,3}=1\pm 2i\]

OpenStudy (anonymous):

but how to find the first step?

OpenStudy (anonymous):

well i see - something you have to get used to ^^ thanks a lot

OpenStudy (anonymous):

Neasd, you can use the 'rational root theorem' to test for rational roots.

OpenStudy (anonymous):

The roots will be of the form\[\frac{p}{q}=\frac{\pm \left\{ factors.of.constant.term \right\}}{\left\{ factors.of.coefficient .of.highest.power \right\}}\]

OpenStudy (anonymous):

So here,\[\frac{p}{q}=\frac{\pm \left\{ 1,2,5,10 \right\}}{\left\{ 1 \right\}}\]

OpenStudy (anonymous):

\[x=\frac{p}{q}=\frac{-2}{1}=-2\]is one element of the set. When you test this, you find\[(-2)^3+(-2)+10=-8-2+10=0\]so x=-2 is a root. You can then factor \[(x-(-2))=(x+2)\]out of your polynomial.

OpenStudy (anonymous):

Do you know how to use long division on polynomials? That is the next step.

OpenStudy (anonymous):

long division? hm never heard of that

OpenStudy (anonymous):

The aim is to find Q(x) such that\[x^3+x+10=(x+2)Q(x)\]

OpenStudy (anonymous):

It's a little difficult to explain on this site. I'll see if I can find a clip that will show you.

OpenStudy (anonymous):

a i see polynomdivision ;D my first language is german ^^

OpenStudy (anonymous):

yeah i know to do that :D thank you!!

OpenStudy (anonymous):

ah, okay :)

OpenStudy (anonymous):

The division will find the Q(x), which is the quadratic Nikvist found.

OpenStudy (anonymous):

kk thx lucas :)

OpenStudy (anonymous):

You're welcome, Markus :)

OpenStudy (anonymous):

Just want to add, if you test all the possible rational combinations and none of them satisfy the condition of a root (i.e. your polynomial does not go to zero), the polynomial has *no* rational roots (you're then left with irrational or complex).

OpenStudy (anonymous):

ok now i'm prepared :D thx again ;)

OpenStudy (anonymous):

np ;)

OpenStudy (anonymous):

may i ask you, what you are doing atm? student? postdoc etc? :D